OpenStudy (anonymous):

How do you find the common factor for 4,24?

OpenStudy (sandra):

well, a common factor is something that evenly divides into both numbers

OpenStudy (sandra):

so I would start with the biggest - which would be one of the two numbers

OpenStudy (sandra):

does 24 divide into both 4 and 24?

OpenStudy (sandra):

no, it's bigger than 4

OpenStudy (sandra):

does 4 divide into both 4 and 24?

OpenStudy (sandra):

yes! 4/4 = 1, and 24/4 = 6

OpenStudy (sandra):

so 4 is a common factor

OpenStudy (sandra):

how about 3?

OpenStudy (sandra):

nope, 3 does not divide evenly into 4

OpenStudy (sandra):

how about 2? yep, 4/2 = 2, 24/2 = 12

OpenStudy (sandra):

how about 1? well, 1 divides into anything, 4/1 = 4, and 24/1 = 24. so the common factors are 4, 2, and 1